OpenStudy (anonymous):

A juggler throws a ball from an initial height of 4 feet with an initial vertical velocity of 30 feet per second. The height h (in feet) of the ball can be modeled by h = –16t2 + vt + s, where t is the time (in seconds) the ball has been in the air, v is the initial velocity (in feet per second), and s is the initial height.

OpenStudy (anonymous):

Write an equation that gives the height (in feet) of the ball as a function of the time (in seconds) since it left the juggler’s hand.

OpenStudy (anonymous):

If the juggler misses the ball, after how many seconds does it hit the ground?

hero (hero):

hint: s = 4, v = 30

OpenStudy (anonymous):

2/15 Sec ?

hero (hero):

2nd hint: The height will be zero when it hits the ground

hero (hero):

Let me see if that is correct

OpenStudy (anonymous):

Okay.

hero (hero):

I get t = 2 secs

hero (hero):

I will post the full steps

OpenStudy (anonymous):

I was thinking 4, lol. Looks like i was off by 2.

OpenStudy (anonymous):

Wanna join in @eyad ??

hero (hero):

h = -16t^2 + 30t + 4 0 = -16t^2 + 30t + 4 16t^2 - 30t - 4 = 0 2(8t^2 - 15t - 2) = 0 8t^2 - 15t - 2 = 0 8t^2 - 16t + t - 2 = 0 8t(t-2)+1(t-2) = 0 (t - 2)(8t + 1) = 0 t - 2 = 0 8t + 1 = 0 t = 2 only

OpenStudy (anonymous):

Yup . It's 2.

hero (hero):

How do you know?

hero (hero):

Is the answer in the back of the book?

OpenStudy (anonymous):

Nope . But i based it on an example that was in my math book.

hero (hero):

standby, I have to verify this answer.

hero (hero):

I just verified that my setup is correct, so that means the answer is also correct.

OpenStudy (anonymous):

Yay!

hero (hero):

I see. What chapter are you on?

OpenStudy (anonymous):

9.. About to be on 10.

hero (hero):

Nice. Making huge progress I see.

OpenStudy (anonymous):

Yup !!! :) I'm ready to finish .

hero (hero):

I know you are :D

OpenStudy (anonymous):

I gotta finish some homework. Ttyl